# When Robustness Meets Conservativeness: Conformalized Uncertainty Calibration for Balanced Decision Making

### v1

@@ Abstract
Robust optimization safeguards decisions against uncertainty by optimizing against worst-case scenarios, yet their effectiveness hinges on a prespecified robustness level that is often chosen ad hoc, leading to either insufficient protection or overly conservative and costly solutions. Recent approaches using conformal prediction construct data-driven uncertainty sets with finite-sample coverage guarantees, but they still fix coverage targets a priori and offer little guidance for selecting robustness levels. We propose a new framework that provides distribution-free, finite-sample guarantees on both miscoverage and regret for any family of robust predict-then-optimize policies. Our method constructs valid estimators that trace out the miscoverage–regret Pareto frontier, enabling decision-makers to reliably evaluate and calibrate robustness levels according to their cost–risk preferences. The framework is simple to implement, broadly applicable across classical optimization formulations, and achieves sharper finite-sample performance than existing approaches. These results offer the first principled data-driven methodology for guiding robustness selection and empower practitioners to balance robustness and conservativeness in high-stakes decision-making.
@@ 1 Introduction
Decision-making under uncertainty is a fundamental challenge in operations research, machine learning, and economics. From utilities preparing for extreme weather [Chen et al., 2025] to financial institutions managing investment risk [Fabozzi et al., 2007], decision-makers must act without knowing the exact outcomes of their choices. Robust optimization (RO) has emerged as a leading paradigm for addressing this challenge: rather than optimizing for the most likely outcome, RO seeks decisions that remain effective under the worst-case realization within a prespecified uncertainty set [Beyer and Sendhoff, 2007]. For instance, grid operators may schedule generator dispatch robust to worst-case demand fluctuations [Bertsimas et al., 2012], while portfolio managers may hedge against worst-case returns consistent with historical data [Goldfarb and Iyengar, 2003].
@@
A standard way to operationalize the RO problems is through the predict-then-optimize (PTO) framework [Bertsimas and Kallus, 2020]: a predictive model first suggests an uncertainty set for the outcome, and the decision-maker then chooses a decision that minimizes the worst-case loss within that set. In traditional RO, this uncertainty set is pre-specified, often as a polyhedral or ellipsoidal region around forecasts [Gabrel et al., 2014]. Its size is governed by a robustness parameter, typically selected through heuristics. For instance, setting the radius to a multiple of the forecast variance. More recently, conformal prediction (CP) has emerged as a data-driven alternative for defining uncertainty sets, providing finite-sample coverage guarantees that the realized outcome lies within the set with a prescribed probability [Vovk et al., 2005]. This approach has also been incorporated into robust optimization, where CP-based sets define the uncertainty region used for decision-making [Patel et al., 2024b].
@@
While these approaches have expanded the scope and rigor of robust decision-making, they leave a critical question unresolved: how should one choose the robustness level? In RO, the radius of the uncertainty set—or in CP, the confidence level—directly governs the trade-off between protection and efficiency. In practice, this choice is typically made heuristically: by fixing conventional confidence levels (e.g., 95%) or by tuning against historical simulations [Lam and Zhou, 2017]. Such methods provide little theoretical guidance, and when data are scarce, simulation-based estimates can be unreliable. This gap makes it difficult for practitioners to balance safety with performance, and decisions are often more conservative than necessary.
@@
In this paper, we propose a new statistical framework that enables decision-makers to evaluate and select robustness levels with explicit trade-offs. Our method is inspired by conformal risk control (CRC) [Angelopoulos et al., 2022], but inverts its logic. Standard CRC begins by prescribing a target risk level and adaptively constructs a set or decision rule to satisfy that risk constraint. By contrast, our inverse CRC framework takes an uncertainty set as given, and then certifies the actual risk level it incurs. This inversion is crucial: it allows us to quantify both miscoverage (the probability that the outcome falls outside the set) and regret (the expected cost relative to an oracle) across all risk levels with rigorous guarantees. Practically, this enables decision-makers to explore the trade-off between robustness and conservativeness in a reliable manner. For example, in the stylized setting of Figure 1, our method may show that reducing coverage by up to 2525% can yield a cost reduction of at least \$$150$. Such insights enable decision-makers to adjust robustness levels adaptively, achieving more cost-effective solutions rather than relying on ad hoc or overly conservative heuristics.
@@
The approach is also straightforward to implement: given calibration data, we construct valid estimators for miscoverage and regret across all robustness levels and use them to trace out a Pareto frontier of risk trade-offs under a mild condition. Theoretical results show that these estimators are reliable, enjoy sharp finite-sample guarantees, and apply broadly to various robust optimization problems. Our numerical results demonstrate that the estimator successfully traces Pareto frontiers that are both valid and tight, and can also effectively guide decision-makers toward robustness parameters that are near-optimal for achieving the desired miscoverage–regret tradeoff.
@@
Our work builds on the broad literature on PTO problems, where a decision-maker observes covariates and must act before outcomes are realized [Bertsimas and Kallus, 2020]. The PTO paradigm has been studied both in two-stage approaches, where one first learns a predictive model and then optimizes with respect to the prediction, and in decision-focused methods that train predictors end-to-end with the downstream objective in mind [Mandi et al., 2024]. Our work complements both streams: instead of introducing new predictors or losses, we develop finite-sample guarantees for evaluating the risk of any robust PTO policy family, enabling practitioners to quantify robustness–conservativeness trade-offs without altering their learning pipeline.
Within PTO, robust optimization (RO) and distributionally robust optimization (DRO) are widely used to hedge against uncertainty. RO approaches define a deterministic uncertainty set and minimize worst-case cost [Ben-Tal and Nemirovski, 2002, Beyer and Sendhoff, 2007, Gabrel et al., 2014], while DRO places ambiguity sets around distributions, typically $\phi$-divergence or Wasserstein balls, and exploits convex duality to reformulate the robust problem [Rahimian and Mehrotra, 2019, Lin et al., 2022]. These methods yield increasingly conservative selectors as the robustness parameter grows. Our framework is orthogonal: given any nested family of uncertainty sets, whether from RO or DRO, we provide distribution-free bounds on realized risks.
A related line of research leverages conformal prediction (CP) to construct uncertainty sets for robust optimization [Patel et al., 2024a, Patel et al., 2024b, Chen et al., 2025]. Here, CP provides data-driven prediction set with finite-sample marginal coverage guarantees [Papadopoulos et al., 2002, Vovk et al., 2005, Zhang et al., 2025], and the robust decision is optimized against these sets. While effective, this approach prescribes the coverage target a priori and tunes the set to achieve it [Patel et al., 2024b]. In contrast, our method inverts the perspective: we take prediction set as given—whether from CP, RO, or DRO—and quantify the actual risks it incurs [Gauthier et al., 2025a, Zhou et al., 2025]. This inversion avoids adaptivity penalties, provides simultaneous guarantees for miscoverage and regret, and ties the risk trade-off curve to the Pareto frontier under majorant consistency.
Our framework is related to the multi-objective optimization literature [Marler and Arora, 2004, Gunantara, 2018], as both aim to characterize the Pareto frontier that formalizes trade-offs among competing goals. The key distinction is that we avoid the nested structure often required in multi-objective formulations: instead of embedding one optimization problem (e.g., robust optimization) inside another (e.g., regret minimization), we exploit the simple structure of the miscoverage–regret trade-off. By directly enumerating over robustness parameters and certifying risk trade-offs from calibration data via conformal techniques, our approach yields a lightweight procedure that admits explicit finite-sample guarantees—something generally unattainable with multi-objective optimization frameworks.
Finally, our work is related to conformal risk control (CRC) [Angelopoulos et al., 2022] and its extensions such as [Farinhas et al., 2023, Teneggi et al., 2023]. Standard CRC prescribes a risk target and ensures validity, while backward or inverse conformal methods start from candidate sets and estimate their realized risk, sometimes via e-values [Balinsky and Balinsky, 2024, Gauthier et al., 2025b]. We differ in three ways: ($i$) we analyze fixed a priori robustness levels, allowing tighter analytic bounds; ($ii$) we extend beyond coverage to regret of a robust decision relative to the oracle; and (iiiiii) our estimator attains validity with sharper finite-sample error rates than other methods. In this way, our work unifies PTO, RO, and CP, but with a new emphasis on jointly certifying robustness and conservativeness along the Pareto frontier.
@@ 2 Problem Setup
We consider a general decision-making problem under uncertainty [Bertsimas and Kallus, 2020]. A decision-maker observes a contextual feature vector $X\in\mathcal{X}$ and must choose a decision variable $z\in\mathcal{Z}$ before the outcome $Y\in\mathcal{Y}$ is realized. The pair $(X,Y)$ is drawn from an underlying distribution $\mathcal{P}$. The quality of a decision $z$ in the presence of outcome $y$ is measured by a cost function $f:\mathcal{Y}\times\mathcal{Z}\to\mathbb{R}$.
@@
$\Big\{\mathcal{C}_{\lambda}(X)\subseteq\mathcal{Y}:\lambda\in\Lambda\subseteq\mathbb{R}_{+}\Big\},$
indexed by a robustness parameter $\lambda$, with the convention that larger $\lambda$ yields larger sets: $\lambda_{1}\leq\lambda_{2}\ \Rightarrow\ \mathcal{C}_{\lambda_{1}}(X)\subseteq\mathcal{C}_{\lambda_{2}}(X)$ for all $X$. For instance, one may construct $\mathcal{C}_{\lambda}(X)$ as an euclidean ball centered at the conditional mean $\hat{g}(X)\approx\mathbb{E}[Y|X]$, where $\hat{g}$ is estimated from data, and take $\lambda$ as the radius of the ball. Given $X$, the robust decision is obtained by solving:
$z_{\lambda}^{*}(X)\coloneqq\arg\min_{z\in\mathcal{Z}}\max_{y\in\mathcal{C}_{\lambda}(X)}f(y,z).$ (1)
Figure 2: Illustration of the problem setup with a simple robust linear optimization problem. Panel (a) shows the miscoverage-regret tradeoff across three robustness parameters ($\lambda_{1},\lambda_{2},\lambda_{3}$). Panel (b) depicts the corresponding robust decisions ($z_{1},z_{2},z_{3}$), which differ from the expected optimal decision $z^{*}$. Panel (c) illustrates an $\ell_{\infty}$-ball uncertainty set $\mathcal{C}_{\lambda}$ for the outcome variable $Y$, with the three adversarial outcome vectors ($y_{1},y_{2},y_{3}$) labeled. The robustness parameters correspond to the radii of these $\ell_{\infty}$-norm sets.
A key challenge is to select $\lambda$ so that the robust decision strikes the right balance between regret and miscoverage as illustrated by Figure 2: If the uncertainty set $\mathcal{C}_{\lambda}(X)$ is too small, the decision may fail to provide sufficient protection against adverse outcomes; If the set is too large, the solution to (1) may be overly conservative and incur an undesirably large regret. We assess each $\lambda$ using two performance metrics given a pair of $(X,Y)$:
$I_{\lambda}(X,Y)$ $\coloneqq\mathbbm{1}\left[Y\notin\mathcal{C}_{\lambda}(X)\right],$
@@
Specifically, for a given pair $(X,Y)$, $I_{\lambda}\in\{0,1\}$ indicates whether the realized outcome $Y$ is excluded from the uncertainty set $\mathcal{C}_{\lambda}(X)$, and $R_{\lambda}$ measures the performance gap of adopting robust decision $z_{\lambda}^{*}(X)$ relative to the oracle decision under the realized $Y$.
@@
$\hat{\alpha}_{I}(\lambda)\geq$ $~\mathbb{E}\left[I_{\lambda}(X,Y)\right]=1-\mathbb{P}\left[Y\in\mathcal{C}_{\lambda}(X)\right],$ (2)
@@ 3 Methodology
This section first introduces a general statistical framework for “inverse” conformal risk control motivated by [Angelopoulos et al., 2022, Patel et al., 2024b], designed to estimate an upper bound on risk defined with respect to an arbitrary loss function. We then demonstrate how this framework can be applied to estimate both miscoverage rate (2) and regret (3) of a robust solution to problem (1). Finally, we show that these estimates are reliable and valid, i.e., the expected upper bound is guaranteed to hold, and that they directly characterize the certified Pareto frontier as the robustness level $\lambda$ varies.
@@ 3.1 Proposed Algorithm
Let $\ell_{\lambda}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ be a family of nonnegative loss (e.g., miscoverage or regret), where $\ell_{\lambda}$ is monotone in $\lambda$ according to the definition of $\mathcal{C}_{\lambda}$. We assume access to a calibration dataset of $n$ paired samples, denoted as $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Given a test data pair $(X_{n+1},Y_{n+1})$, our goal is to construct an estimator, denoted $\hat{\alpha}_{\ell}(\lambda)$, from the calibration data $\mathcal{D}$ such that
@@
This objective subsumes both (2) and (3), and can be interpreted as an “inverse” form of the conformal risk control (CRC) [Angelopoulos et al., 2022]. Specifically, CRC begins by prescribing a desired risk level $\alpha$ on the expected loss and then finds $\lambda$ that guarantees:
@@
This assumption is standard in the conformal prediction literature [Vovk et al., 2005] and subsumes many common settings, including the case of i.i.d. sampling. Furthermore, we assume that there exists a known uniform upper bound on the regret.
Assumption 2.
There exists a known constant $B\geq 0$ such that for all $\lambda\in\Lambda$, $\mathbb{P}\left\{0\leq\ell_{\lambda}(X,Y)\leq B\right\}=1.$
2 typically requires some prior knowledge of the loss function to verify. In many cases, however, the bound $B$ is straightforward to obtain—for example, when the loss function is inherently bounded or when the distribution $\mathcal{P}$ has bounded support.
@@
Proposition 1.
Let $\bar{\ell}_{n}(\lambda)$ denote the average calibration loss, i.e., $\bar{\ell}_{n}(\lambda)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell_{\lambda}(x_{i},y_{i}).$ The proposed risk estimator is defined as
$\hat{\alpha}_{\ell}(\lambda)=\inf\Bigg\{$ $\alpha\in(0,B):$ (5)
$B-\bar{\ell}_{n}(\lambda)\geq\frac{\lceil(n+1)(B-\alpha)\rceil}{n}\Bigg\},$
In practice, a more convenient approximation is
@@
which can be derived by relaxing the ceiling operator in (5). This form requires only the empirical average calibration loss $\bar{\ell}_{n}(\lambda)$ plus an adjustment $B/(n+1)$ that accounts for the worst-case contribution of the unseen test point, guaranteed to lie in $[0,B]$ by Assumption 2. This conservative correction is crucial for ensuring validity: without it, the empirical estimate alone would not guarantee coverage on new data. The complexity of computing $\bar{\ell}_{n}(\lambda)$ is linear in the number of calibration samples, and the adjustment is constant, making $\hat{\alpha}(\lambda)$ efficient to evaluate for each candidate $\lambda$.
Finally, we construct the miscoverage–regret frontier using the above estimator as follows: For each candidate robustness level $\lambda$, we first solve the robust problem in (1) to obtain the decision rule $z_{\lambda}^{*}(\cdot)$, then evaluate the induced miscoverage and regret losses on calibration data. These empirical averages are corrected using the conservative adjustment in Proposition 1 to yield valid upper-bound estimates of risk. Collecting the pairs $\big(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda)\big)$ across all $\lambda$ forms the empirical frontier, which is then Pareto-pruned to remove dominated points. The pseudocode in Algorithm 1 summarizes the entire proposed method.
Algorithm 1 CREME: Conformal REgret-Miscoverage Estimate
[…]
2:for $\lambda\in\{\lambda_{i}\}$ do
3:  $\bar{I}_{n}(\lambda)\leftarrow\sum_{i=1}^{n}I_{\lambda}(x_{i},y_{i})/n$;
4:  $\bar{R}_{n}(\lambda)\leftarrow\sum_{i=1}^{n}R_{\lambda}(x_{i},y_{i})/n$;
5:  $\tilde{\alpha}_{R}(\lambda)\leftarrow$ compute (6) with $\bar{R}_{n}(\lambda)$, $B=R_{\text{max}}$;
6:  $\tilde{\alpha}_{I}(\lambda)\leftarrow$ compute (6) with $\bar{I}_{n}(\lambda)$, $B=1$;
[…]
8:end for
9: Remove dominated points from $\hat{\mathcal{F}}$ (keep only tuples not weakly worse in both coordinates);
10:return Certified Pareto frontier $\hat{\mathcal{F}}$.

### v2

# Calibrating Decision Robustness via Inverse Conformal Risk Control

@@ Abstract
Robust optimization safeguards decisions against uncertainty by optimizing against worst-case scenarios, yet their effectiveness hinges on a prespecified robustness level that is often chosen ad hoc, leading to either insufficient protection or overly conservative and costly solutions. Recent approaches using conformal prediction construct data-driven uncertainty sets with finite-sample coverage guarantees, but they still fix coverage targets a priori and offer little guidance for selecting robustness levels. We propose a new framework that provides distribution-free, finite-sample guarantees on both miscoverage and regret for any family of robust predict-then-optimize policies. Our method constructs valid estimators that trace out the miscoverage–regret Pareto frontier, enabling decision-makers to reliably evaluate and calibrate robustness levels according to their cost–risk preferences. The framework is simple to implement, broadly applicable across classical optimization formulations, and achieves sharper finite-sample performance. This paper offers a principled data-driven methodology for guiding robustness selection and empowers practitioners to balance robustness and conservativeness in high-stakes decision-making.
Conformal prediction, robust optimization, multiobjective optimization
@@ 1 Introduction
Decision-making under uncertainty is a fundamental challenge in operations research, machine learning, and economics. From utilities preparing for extreme weather (Chen et al., 2025) to financial institutions managing investment risk (Fabozzi et al., 2007), decision-makers must act without knowing the exact outcomes of their choices. Robust optimization (RO) has emerged as a leading paradigm for addressing this challenge: rather than optimizing for the most likely outcome, RO seeks decisions that remain effective under the worst-case realization within a prespecified uncertainty set (Beyer and Sendhoff, 2007). For instance, grid operators may schedule generator dispatch robust to worst-case demand fluctuations (Bertsimas et al., 2012), while portfolio managers may hedge against worst-case returns consistent with historical data (Goldfarb and Iyengar, 2003).
@@
A standard way to operationalize the RO problems is through the predict-then-optimize (PTO) framework (Bertsimas and Kallus, 2020): a predictive model first suggests an uncertainty set for the outcome, and the decision-maker then chooses a decision that minimizes the worst-case loss within that set. In traditional RO, this uncertainty set is pre-specified, often as a polyhedral or ellipsoidal region around forecasts (Gabrel et al., 2014). Its size is governed by a robustness parameter, typically selected through heuristics. For instance, setting the radius to a multiple of the forecast variance. More recently, conformal prediction (CP) has emerged as a data-driven alternative for defining uncertainty sets, providing finite-sample coverage guarantees that the realized outcome lies within the set with a prescribed probability (Vovk et al., 2005). This approach has also been incorporated into robust optimization, where CP-based sets define the uncertainty region used for decision-making (Patel et al., 2024a).
@@
While these approaches have expanded the scope and rigor of robust decision-making, they leave a critical question unresolved: how should one choose the robustness level? In RO, the radius of the uncertainty set—or in CP, the confidence level—directly governs the trade-off between protection and efficiency. In practice, this choice is typically made heuristically: by fixing conventional confidence levels (e.g., 95%) or by tuning against historical simulations (Lam and Zhou, 2017). Such methods provide little theoretical guidance, and when data are scarce, simulation-based estimates can be unreliable. This gap makes it difficult for practitioners to balance safety with performance, and decisions are often more conservative than necessary.
@@
In this paper, we propose a new statistical framework that enables decision-makers to evaluate and select robustness levels with explicit trade-offs. Our method is inspired by conformal risk control (CRC) (Angelopoulos et al., 2022), but inverts its logic. Standard CRC begins by prescribing a target risk level and adaptively constructs a set or decision rule to satisfy that risk constraint. By contrast, our inverse CRC framework takes an uncertainty set as given, and then certifies the actual risk level it incurs. This inversion is crucial: it allows us to quantify both miscoverage (the probability that the outcome falls outside the set) and regret (the expected cost relative to an oracle) across all risk levels with rigorous guarantees. Practically, this enables decision-makers to explore the trade-off between robustness and conservativeness in a reliable manner. For example, in the stylized setting of Figure 1, our method may show that reducing coverage by up to 2525% can yield a cost reduction of at least \$$150$. Such insights enable decision-makers to adjust robustness levels adaptively, achieving more cost-effective solutions rather than relying on ad hoc or overly conservative heuristics.
@@
The approach is also straightforward to implement: given calibration data, we construct valid estimators for miscoverage and regret across all robustness levels and use them to trace out a Pareto frontier of risk trade-offs under a mild condition. If the robustness level is chosen post hoc by a human decision-maker based on the estimated trade-offs, we explicitly account for this selection by performing a final recalibration on a split dataset, which restores finite-sample validity for the reported miscoverage and regret guarantees. Theoretical results show that these estimators are reliable, enjoy sharp finite-sample guarantees, and apply broadly to various robust optimization problems. Our numerical results demonstrate that the estimator successfully traces Pareto frontiers that are both valid and tight, and can also effectively guide decision-makers toward robustness parameters that are near-optimal for achieving the desired miscoverage–regret tradeoff.
@@
Our work is related to the predict-then-optimize (PTO) literature, which studies decision-making under uncertainty via two-stage pipelines or decision-focused learning (Bertsimas and Kallus, 2020; Mandi et al., 2024). Within this paradigm, robust optimization (RO) and distributionally robust optimization (DRO) address uncertainty by optimizing against deterministic or distributional ambiguity sets (Ben-Tal and Nemirovski, 2002; Beyer and Sendhoff, 2007; Gabrel et al., 2014; Rahimian and Mehrotra, 2019; Lin et al., 2022). Complementary to these frameworks, we adopt a diagnostic viewpoint: rather than prescribing a particular robust decision, our focus is on characterizing and certifying the risk trade-offs induced by different robustness levels, thereby supporting decision-making through principled auditing and calibration. This viewpoint also connects our work to Pareto-front analysis in multi-objective optimization (Marler and Arora, 2004; Gunantara, 2018), where trade-offs are typically explored via parameter enumeration. Building on this literature, we construct a certified Pareto front with explicit finite-sample guarantees using conformal prediction—an element that, to our knowledge, has not yet been incorporated into Pareto-front analysis.
Conformal prediction (CP) provides a general framework for uncertainty quantification with finite-sample validity guarantees (Papadopoulos et al., 2002; Vovk et al., 2005). Recent work has focused on improving the efficiency of CP in multi-dimensional settings by adapting prediction sets to richer geometric structures, including ellipsoidal sets (Messoudi et al., 2022; Xu et al., 2024), convex template-based sets (Tumu et al., 2024), and density-adaptive constructions (Izbicki et al., 2022; Wang et al., 2023; Messoudi et al., 2021; Sun and Yu, 2023). These contributions have laid important groundwork for a growing body of research integrating CP with robust optimization (RO), where CP is used to construct data-driven uncertainty sets for decision-making (Johnstone and Cox, 2021). Representative recent works include the use of probabilistic CP in predict-then-optimize pipelines (Patel et al., 2024a), CP-based robust control for linear–quadratic regulator systems (Patel et al., 2024b), decision-focused frameworks that leverage CP to guarantee robustness (Cortes-Gomez et al., 2024), and multi-layer robust optimization formulations for grid operations informed by CP (Chen et al., 2025). Our work is complementary to this line of research. Rather than using CP to construct uncertainty set, we consider settings in which the uncertainty set family is specified a priori, with CP used to certify risk trade-offs as the set size varies.
Conformal risk control (CRC) generalizes conformal prediction to construct loss functions with prespecified miscoverage guarantee (Angelopoulos et al., 2022). Subsequent work has substantially extended the CRC framework in multiple directions: (Farinhas et al., 2023) studies CRC under non-exchangeable data; (Blot et al., 2025) proposes an adaptive CRC procedure inspired by adaptive conformal prediction (Zaffran et al., 2022) to better exploit conditional information; (Teneggi et al., 2023) extends CRC to multidimensional outputs with applications to image generation; (Cohen et al., 2024) introduces a cross-validation-based approach to mitigate data scarcity; (Yeh et al., 2025) generalizes CRC to optimization problems with control over tail-sensitive risks; and (Luo and Zhou, 2025) applies CRC to image segmentation tasks. In contrast, our work addresses the inverse problem of CRC: given a fixed loss function, how can we estimate its miscoverage? This inverse perspective can be viewed as a generalization of inverse conformal prediction (Prinster et al., 2022; Singh et al., 2024; Zhou et al., 2025; Gauthier et al., 2025a) to CRC.
Finally, our work relates to post-hoc selection bias arising from data-dependent parameter selection (Cox, 1975). Recent works have proposed stability-based approaches (Zrnic and Jordan, 2023; Hegazy et al., 2025) and e-value-based methods in CP settings (Balinsky and Balinsky, 2024; Gauthier et al., 2025b, a). In contrast, our work follows a more typical data-splitting strategy (Cox, 1975; Wasserman and Roeder, 2009; Fithian et al., 2014; Rasines and Young, 2023).
@@ 2 Problem Setup
Figure 2: Illustration of the problem setup with a simple robust linear optimization problem. Panel (a) shows the miscoverage-regret tradeoff across three robustness parameters ($\lambda_{1},\lambda_{2},\lambda_{3}$). Panel (b) depicts the corresponding robust decisions ($z_{1},z_{2},z_{3}$), which differ from the expected optimal decision $z^{*}$. Panel (c) illustrates an $\ell_{\infty}$-ball uncertainty set $\mathcal{U}_{\lambda}$ for the outcome variable $Y$, with the three adversarial outcome vectors ($y_{1},y_{2},y_{3}$) labeled. The robustness parameters correspond to the radii of these $\ell_{\infty}$-norm sets.
We consider a general decision-making problem under uncertainty (Bertsimas and Kallus, 2020). A decision-maker observes a contextual feature vector $X\in\mathcal{X}$ and must choose a decision variable $z\in\mathcal{Z}$ before the outcome $Y\in\mathcal{Y}$ is realized. The pair $(X,Y)$ is drawn from an underlying distribution $\mathcal{P}$. The quality of a decision $z$ in the presence of outcome $y$ is measured by a cost function $f:\mathcal{Y}\times\mathcal{Z}\to\mathbb{R}$.
@@
$\Big\{\mathcal{U}_{\lambda}(X)\subseteq\mathcal{Y}:\lambda\in\Lambda\subseteq\mathbb{R}_{+}\Big\},$
indexed by a robustness parameter $\lambda$, with the convention that larger $\lambda$ yields larger sets: $\lambda_{1}\leq\lambda_{2}\ \Rightarrow\ \mathcal{U}_{\lambda_{1}}(X)\subseteq\mathcal{U}_{\lambda_{2}}(X)$ for all $X$. For instance, one may construct $\mathcal{U}_{\lambda}(X)$ as an euclidean ball centered at the conditional mean $\hat{g}(X)\approx\mathbb{E}[Y|X]$, where $\hat{g}$ is estimated from data, and take $\lambda$ as the radius of the ball. Given $X$, the robust decision is obtained by solving:
$z_{\lambda}^{*}(X)\coloneqq\arg\min_{z\in\mathcal{Z}}\max_{y\in\mathcal{U}_{\lambda}(X)}f(y,z).$ (1)
A key challenge is to select $\lambda$ so that the robust decision strikes the right balance between regret and miscoverage as illustrated by Figure 2: If the uncertainty set $\mathcal{U}_{\lambda}(X)$ is too small, the decision may fail to provide sufficient protection against adverse outcomes; If the set is too large, the solution to (1) may be overly conservative and incur an undesirably large regret. We assess each $\lambda$ using two performance metrics given a pair of $(X,Y)$:
$I_{\lambda}(X,Y)$ $\coloneqq\mathbbm{1}\left[Y\notin\mathcal{U}_{\lambda}(X)\right],$
@@
Specifically, for a given pair $(X,Y)$, $I_{\lambda}\in\{0,1\}$ indicates whether the realized outcome $Y$ is excluded from the uncertainty set $\mathcal{U}_{\lambda}(X)$, and $R_{\lambda}$ measures the performance gap of adopting robust decision $z_{\lambda}^{*}(X)$ relative to the oracle decision under the realized $Y$.
@@
$\hat{\alpha}_{I}(\lambda)\geq$ $~\mathbb{E}\left[I_{\lambda}(X,Y)\right]=1-\mathbb{P}\left[Y\in\mathcal{U}_{\lambda}(X)\right],$ (2)
@@ 3 Methodology
This section first introduces a general statistical framework for “inverse” conformal risk control motivated by (Angelopoulos et al., 2022; Patel et al., 2024a), designed to estimate an upper bound on risk defined with respect to arbitrary loss functions, and construct a Pareto frontier over both miscoverage rate (2) and regret (3) of a robust solution to problem (1). Then, we show that these estimates are reliable and valid, i.e., the expected upper bound is guaranteed to hold, and that they directly characterize the certified Pareto frontier as the robustness level $\lambda$ varies. Finally, we extend our algorithm to retain validity when used for recalibration after selection by adopting a data splitting technique.
@@ 3.1 Proposed Algorithm
For any prespecified $\lambda\in\Lambda$, let $\ell_{\lambda}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ be a family of nonnegative loss (e.g., miscoverage or regret), where $\ell_{\lambda}$ is monotone in $\lambda$ according to the definition of $\mathcal{U}_{\lambda}$. We assume access to a calibration dataset of $n$ paired samples, denoted as $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Given a test data pair $(X_{n+1},Y_{n+1})$, our goal is to construct an estimator, denoted $\hat{\alpha}_{\ell}(\lambda)$, from the calibration data $\mathcal{D}$ such that
@@
This objective subsumes both (2) and (3), and can be interpreted as an “inverse” form of the conformal risk control (CRC) (Angelopoulos et al., 2022). Specifically, CRC begins by prescribing a desired risk level $\alpha$ on the expected loss and then finds $\lambda$ that guarantees:
@@
This assumption is standard in the conformal prediction literature (Vovk et al., 2005) and subsumes many common settings, including the case of i.i.d. sampling. Following (Angelopoulos et al., 2022), we assume that there exists the essential supremum of the regret is bounded.
Assumption 3.2.
There exists a constant $B\geq 0$ such that, for all $\lambda\in\Lambda$,
$\ell_{\lambda}(X,Y)\in[0,B]\quad\text{a.s.}$
The bound $B$ is known or can be derived from prior knowledge of the loss function, e.g., for miscoverage losses $I_{\lambda}\in\{0,1\}$, one may take $B=1$; more generally, $B$ can be computed analytically when data has bounded support.
@@
Definition 3.3.
Let $\bar{\ell}_{n}(\lambda)$ denote the average calibration loss, i.e., $\bar{\ell}_{n}(\lambda)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell_{\lambda}(x_{i},y_{i}).$ The proposed risk estimator $\hat{\alpha}_{\ell}(\lambda)$ is defined as
$\inf_{\alpha\in(0,B)}\Bigg\{\alpha:B-\bar{\ell}_{n}(\lambda)\geq\frac{\lceil(n+1)(B-\alpha)\rceil}{n}\Bigg\},$ (5)
In practice, a more convenient approximation to (5) is
@@
which can be derived by relaxing the ceiling operator in (5). This form requires only the empirical average calibration loss $\bar{\ell}_{n}(\lambda)$ plus an adjustment $B/(n+1)$ that accounts for the worst-case contribution of the unseen test point, guaranteed to lie in $[0,B]$ by Assumption 3.2. This conservative correction is crucial for ensuring validity. Otherwise, the empirical estimate alone would not guarantee coverage on new data. The complexity of computing $\bar{\ell}_{n}(\lambda)$ is linear in the number of calibration samples, and the adjustment is constant, making $\hat{\alpha}(\lambda)$ itself efficient to evaluate.
Finally, we construct the miscoverage-regret frontier using the above estimator as follows: for each candidate robustness level $\lambda\in\Lambda$, we first solve the robust problem in (1) to obtain the decision rule $z_{\lambda}^{*}(\cdot)$, then evaluate the induced miscoverage and regret losses on calibration data. These empirical averages are corrected using the conservative adjustment in Proposition 3.3 to yield valid upper-bound estimates of risk. Collecting the pairs $\big(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda)\big)$ across all $\lambda$ forms the empirical frontier, which is then Pareto-pruned to remove dominated points.
We note that the primary computational bottleneck of the enumeration-based procedure is the need to solve (1) for each (discretized) element of the index set $\Lambda$. However, this cost is often manageable when $z^{\ast}_{\lambda}(\cdot)$ admits closed-form solutions or can be efficiently solved by off-the-shelf solvers, so this computational complexity is standard in the multi-objective optimization literature (Miettinen, 1999). We also validated this through our empirical analysis (Section 4.3).